OpenStudy (anonymous):

Solve the following system of equations: x - 2y = 14 x + 3y = 9

OpenStudy (anonymous):

just subtract them ...and solve for y then the value of y put in any of this eq you will get the reslut..

OpenStudy (anonymous):

get it..?

OpenStudy (anonymous):

Not really.

OpenStudy (geekfromthefutur):

https://www.mathway.com Try that darlin

OpenStudy (anonymous):

You cant set up a system of equations @geekfromthefutur on there....

OpenStudy (geekfromthefutur):

o

OpenStudy (anonymous):

see...it below

OpenStudy (anonymous):

subtracting both eq gives... (x-x)+(-2y-3y)=14-9 => -5y=5 => y=-1 now put y=-1 in any of given eq.you will get.. say eq 1 x-2y=14 =>x-2(-1)=14 =>x=14-2 =>x=12 so sol set is. (x,y)=(12,-1)

OpenStudy (anonymous):

Thank you so much, I'm starting to get it a little better now.

OpenStudy (anonymous):

you welcome..!!